# thanks: A footnote to the article title

### v2

@@ II.4 Achieving Optimal Fractions of Exact Exchange in the r2SCANY@r2SCANX Formulations
II.5 Effects of Exact-exchange Fractions X and Y on the Binding Energy of the Oxygen Molecule
@@
Note that zero-point energy corrections are not included in this analysis and represent a constant, monotonic energy shift of $\sim$0.1 eV/O2.155, 156 The inadequacy of LSDA XC functionals, which drastically overbinds O2, had already been noted by Perdew and Zunger157 in agreement with LSDA-PW92 in Fig. 4a. Similar inaccuracies were also identified for more accurate GGA functionals by Blöchl158 and later by Wang et al.29
@@
II.6 Effects of Exact-Exchange Fractions X and Y on On-site Magnetic Moments
The strong correlation in MiOjs seems to be captured at least in part by density functional approximations for normal correlation through spin symmetry breaking.24, 159, 160, 161, 162 In this work, we have used experimental or nearly experimental magnetic orders (referenced in Table 1).
The artificial delocalization of d electrons due to SIE of XC functionals affects the predicted on-site magnetic moments in transition metals.46, 47, 51, 67, 163, 164, 165 We investigate the variation in error of predicted magnetic moments in our proposed methods r2SCANY@r2SCANX. The magnetic moments are calculated by integrating the net spin density over the projector augmented wave (PAW) potential spheres of the transition metal atoms.
@@ II.6 Effects of Exact-Exchange Fractions X and Y on On-site Magnetic Moments
Finally, we comment on the performance of the meta-GGA LAK in predicting the on-site magnetic moments in MiOjs. Fig. 5c and Fig. 5d show no noticeable improvement of LAK magnetic moments compared to r2SCAN.
II.7 Effects of Exact-exchange Fractions X and Y on the Prediction of Fundamental Band Gaps
Band gaps of materials are typically affected by SIE, with band gaps often underestimated by LSDA, GGA, and meta-GGA functionals.25, 26, 38, 39, 98, 166, 167, 168 We investigate the error in predicted band gaps using our r2SCANY@r2SCANX method. The fundamental band gap is the difference between the lowest unoccupied and the highest occupied orbital energies. Our orbital energies are expectation values of the r2SCANY one-electron generalized Kohn-Sham one-electron Hamiltonian using the r2SCANX orbitals. Y=100% includes the full Fock operator and strongly overestimates the band gaps.
@@ III Discussion
Hybrid XC functionals mix a fraction of exact HF exchange with DFA to improve the accuracy of the electronic structure, thereby correcting functional-driven errors, particularly the SIE. The amount of HF exchange remains a tunable parameter. Strategies proposed to tune the amount of HF exchange in the DFA include: fitting molecular heat of formations to experimental data (Pople’s G1 dataset), for example, the B3PW91 and the B3LYP.66, 169 Appropriate amounts of HF can also be derived from first-principles perturbative theory calculations,69 as is the case for 25% HF exchange in PBE0 and HSE06.70, 71 Galli et al.63 demonstrated that the exact exchange amount can be inferred directly considering the material dielectric constant. Scuseria et al.,72, 73, 74, 75 proposed that the amount of HF can be determined as a function of the electron density at each point in space. We have shown that MiOjs’ oxidation energies generally improve climbing the ladder from LSDA to PBE to r2SCAN, but further improvement is needed.
@@
Addressing Density and Functional Driven Errors Altogether: HF also provides self-interaction-free, albeit uncorrelated, electronic charge densities. Electronic charge densities from HF or hybrid functionals can correct density-driven errors. We have demonstrated that HF (or hybrid functional) electronic charge densities can correct r2SCAN (or its hybrids) total energies for MiOjs by proposing a generalized r2SCANY@r2SCANX method. In r2SCAN@r2SCANX, we mix Yexact HF exchange in the DFA and set the fraction of exact HF exchange used to evaluate the charge density to X%. We have proposed optimal combinations of HF exchange in the charge density and functional definitions, such as r2SCAN10, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, and r2SCAN10@r2SCAN50, which reduce the mean absolute (and relative) errors of predicted electronic and magnetic properties, and oxidation energies of MiOjs.
@@
Kulik and collaborators demonstrated that the ability of DFT+U to localize electrons on transition metals can vary significantly from that of hybrid functionals; these latter tend to localize the minority spin density (of the transition metal) away from the metal, and towards oxygen atoms in transition metal complexes.170, 171 We observe the same behavior in MiOjs. In agreement with Kulik et al.,170, 171 we also show that a higher fraction of HF exchange in the XC functional increases the on-site magnetic moments on transition metal species in MiOjs, which is rationalized by the rise of majority spin in the d-manifolds. Considering these observations, we can qualitatively state that an increasing amount of HF in the DFA increases the ionic character of MiOjs. We quantify these behaviors for Fe2O3 Supplementary Fig. 5 by change in minority- and majority-spin electron numbers. In addition, Supplementary Fig. 6 shows a progressive increase in the electron number on the oxygen sites with increasing X in r2SCANX. This is preceded by a similar increase from LSDA to PBE to r2SCAN. The electron number on the transition-atom site decreases with increasing X. Exceptions to this trend are MiOjs, where strong charge disproportionation of the transition metal is favored, for example, Fe3O4 (Fe3+ and Fe2+), Mn3O4 (Mn3+ and Mn4+). As for the magnetic moments, the electron number on an atom is estimated by integrating the electron density over the PAW sphere.
In addition, while investigating transition metal complexes, Ref. 171 concluded that DFT+U charge densities might differ substantially from hybrid functionals. To test this observation, we compare r2SCAN@r2SCAN+U MiOjs band gaps and oxidation energies (Supplementary Fig. 7) to r2SCAN and r2SCAN10@r2SCAN50 results. We demonstrate that r2SCAN@r2SCAN+U oxidation energies and band gaps are comparable to r2SCAN.
@@
The widely accepted standard for representing uncertainties in experimentally obtained oxidation enthalpies and other thermochemical data requires estimates of the 95% confidence intervals, which all thermochemical tables universally follow.149, 150, 151 Ruscic has proposed that the MAE of data might underestimate the suggested thermochemical uncertainty by a factor as low as 3.5$\times$, thereby invalidating numerous assertions of achieving ”chemical accuracy.”172 For this reason, while benchmarking the accuracy of r2SCANY@r2SCANX, we have also examined the impact of the root mean square error.
@@
Reproducing errorless O2 binding energies is essential for accurate predictions of oxidation energies of MiOjs. LSDA and PBE drastically overbind The O2 molecule with large negative binding energies,29, 157, 158 $\sim$–2.2 eV/O2 and $\sim$–1 eV/O2, respectively, resulting in a systematic shift in predicted oxidation energies. With r2SCAN this negative error drops in magnitude to a sizeable value of $\sim$–0.3 eV/O2. Applying a constant shift (of $\sim$0.3 eV/O2) to correct the spurious O2 overbinding in r2SCAN would further underestimate MiOj oxidation energies (see Fig. 2a), further exacerbating the error in r2SCAN predictions. Previously, similar corrections have been applied to PBE O2 and MiOj oxidation energies.29, 30 Notably, we have demonstrated that r2SCAN10@r2SCAN, r2SCAN10, and r2SCAN10@r2SCAN50 decrease the error in the O2 binding energy to values $\leq$ 0.03 eV/O2 in magnitude, and without requiring any fitting procedures.29, 30
@@
Since the errors of O2 binding energy are minimal with r2SCANY@r2SCANX methods, errors in the oxidation energies of MiOjs must reside in the prediction of the reduced and oxidized oxides. This is a bimodal distribution of r2SCAN oxidation energies (Fig. 2b), consisting of a low error peak for early transition metals, i.e., Ti, V, Cr, reactions, and higher errors for late transition metals, i.e., Mn, Fe, Co reactions. Notably, early transition metals (Ti, V, and Cr) involve a small number of d-electrons in their electronic configurations, probably reducing correlation effects, which become more prominent for late transition metals (Mn, Fe, and Co) with more electrons in d-shells. In addition, redox reactions involving exchanges between transition metals’ closed d-shell configurations to open d-shell configurations or vice versa, such as Ti4+(d0) $\rightarrow$ Ti3+ (d1), or Cu+ (d10) $\rightarrow$ Cu2+ (d9), appear to be in better agreement with experimental data. However, milder correlation effects in closed-shell transition metals make r2SCAN sufficient to describe the electronic structure of their oxides (for example, TiO2, V2O5, etc.).
@@
From these observations, we expect that the electron density by r2SCAN may be too delocalized to capture localization effects in late transition metals, i.e., Mn, Fe, Co, containing a larger number of d electrons. Better electronic charge densities as provided by r2SCAN@r2SCAN50, correct the electron number in the d manifolds, recovering a unimodal distribution (Fig. 2b), with a substantial decrease of oxidation energies errors for MiOjs with M = Mn, Fe, and Co. In addition, due to the functional-driven error, the average error r2SCAN@r2SCAN50 is non-zero, resulting in an offset distribution around the zero-error line. However, in r2SCAN10@r2SCAN50 the functional-driven error reduces, and the unimodal distribution centers near the zero-error line.
@@
Magnetic moments are also typically impacted by SIE and are underestimated due to the overdelocalization of d(f) electrons on transition metal centers by GGA and meta-GGA functionals. As we climb the Jacob’s ladder173 of the XC functionals from LSDA (PW91) to GGA, and to meta-GGAs, the error in magnetic moments decreases accordingly, as observed from the functional MAE distributions of Fig. 5a and Fig. 5b. We observed that the error in magnetic moments drops slowly as the percentage of exact HF exchange in the orbital increases. Compared to r2SCAN, for r2SCAN50, the error in magnetic moment reduces by $\sim$15%, and for HF, the error reduces by $\sim$20%. In contrast, r2SCAN+U minimizes the error in magnetic moments much more mildly, $\sim$5%. Note that the variance in experimentally reported magnetic moments can be as high as +1.00 $\mu_{B}$. For example, magnetic moments in V3+ in V2O3 are reported to vary between 1.2 to 2.37 $\mu_{B}$ (Table 1),121, 122 which corresponds to variations larger than 100%. Despite this considerable uncertainty in experimentally reported magnetic moments, we observed a noticeable decrease in MAE of MiOjs predicted magnetic moments from r2SCAN to r2SCAN50.
@@
Band gaps of MiOjs are typically underestimated by standard DFT predictions, with errors rapidly decreasing when using LSDA (PW91) $\gg$ GGA (PBE) > meta-GGAs (r2SCAN and LAK) > Hybrid meta-GGA (r2SCANX), see Fig. 6c and Fig. 6d. Using r2SCANY@r2SCANX, we demonstrated that errors in band gaps are minimized with approximately 5-10% exact exchange in the functional and the orbital definitions, which differs from the optimal HF fractions minimizing oxidation energy errors of MiOjs. We have shown that band gaps predicted by r2SCAN10@r2SCAN are more accurate than r2SCAN+U, with MAEs of 0.53 eV/O2 and 0.71 eV/O2, respectively.
@@
Materials possess many degrees of freedom embedded in their crystal structures. Deformation of octahedra (and cation polyhedra), octahedral tilting, and Jahn-Teller distortions all decrease the crystal symmetry in materials.24, 161 Additionally, specific magnetic orderings of transition metals in MiOjs, typically determined by neutron scattering and magnetic spectroscopies, further reduce crystal symmetries. On the one hand, standard structural techniques, such as X-ray and neutron diffraction, generally show less sensitivity to such distortions and defects. As such, reported crystal structures of materials may be inherently more symmetric than they appear. On the other hand, band-gap measurements are sensitive to symmetry-breaking motifs, defects in materials, and charge and magnetic orderings. Zunger and collaborators demonstrated that DFT of overly symmetrized MiOjs and other materials tends to close band gaps, predicting “false metals”.24, 159, 160, 161
@@
Starting from the highly symmetrized experimental structures, further optimized with r2SCAN, our band-gap predictions of Ti2O3,101 V2O3,103 and Fe3O4113 appear metallic (Table 1). Upon removing all possible symmetry layers (perturbing atomic positions from high-symmetry sites of large supercell models, imposing ground state magnetic orderings, removing intrinsic symmetry of wavefunctions, orbitals, and time reversal), Ti2O3, V2O3, and Fe3O4 remain metallic. However, the experimentally reported band gaps for Ti2O3, V2O3, and Fe3O4 are smaller than or equal to $\sim$0.2 eV (Table 1), which makes band-gap opening unlikely, even with substantial symmetry breaking as demonstrated by our calculations. We found no lattice symmetry breaking for Ti2O3 and V2O3, while we observed lattice symmetry breaking for Fe3O4. However, there can be multiple symmetry breakings,24, 159 and we may not have found the one with the lowest energy.
Computational Efficiency Considerations of r2SCANY@r2SCANX approaches: It is relevant to comment on the computational costs faced by the r2SCANY@r2SCANX approach. Excluding the computational time required to obtain the initial set of r2SCAN orbitals self-consistently (which are used directly in the r2SCAN10@r2SCAN calculation and as a starting guess for r2SCANX calculation), the non-self-consistent r2SCAN10@r2SCAN calculations are approximately n times faster than the corresponding SCF global hybrid r2SCAN10 calculations. Here, n represents the number of electronic SCF iterations typically required for a global hybrid r2SCANX convergence, generally ranging from 10 to 300, depending on system complexity and total energy convergence criteria.

### v3

@@ II.D Achieving Optimal Fractions of Exact Exchange in the r2SCANY@r2SCANX Formulations
II.E Effects of Exact-exchange Fractions X and Y on the Binding Energy of the Oxygen Molecule
@@
Note that zero-point energy corrections are not included in this analysis and represent a constant, monotonic energy shift of $\sim$0.1 eV/O2.154, 155 The inadequacy of LSDA XC functionals, which drastically overbinds the O2 molecule had already been noted by Perdew and Zunger156 in agreement with LSDA-PW92 in Fig. 4a. Similar inaccuracies were also identified for more accurate GGA functionals by Blöchl157 and later by Wang et al.29
@@
II.F Effects of Exact-Exchange Fractions X and Y on On-site Magnetic Moments
The strong correlation in MiOjs seems to be captured at least in part by density functional approximations for normal correlation through spin symmetry breaking.24, 158, 159, 160, 161 In this work, we have used experimental or nearly experimental magnetic orders (referenced in Table 1).
The artificial delocalization of d electrons due to SIE of XC functionals affects the predicted on-site magnetic moments in transition metals.46, 47, 51, 67, 162, 163, 164 We investigate the variation in error of predicted magnetic moments in our proposed methods r2SCANY@r2SCANX. The magnetic moments are calculated by integrating the net spin density over the projector augmented wave (PAW) potential spheres of the transition metal atoms.
@@ II.F Effects of Exact-Exchange Fractions X and Y on On-site Magnetic Moments
Finally, we comment on the performance of the meta-GGA LAK in predicting the on-site magnetic moments in MiOjs. Fig. 5c and Fig. 5d show an improvement from LAK over r2SCAN by only about 5% for the magnetic moments.
II.G Effects of Exact-exchange Fractions X and Y on the Prediction of Fundamental Band Gaps
Band gaps of materials are typically affected by SIE, with band gaps often underestimated by LSDA, GGA, and meta-GGA functionals.25, 26, 38, 39, 98, 165, 166, 167 We investigate the error in predicted band gaps using our r2SCANY@r2SCANX method. The fundamental band gap is the difference between the lowest unoccupied and the highest occupied orbital energies. Our orbital energies are expectation values of the r2SCANY one-electron generalized Kohn-Sham Hamiltonian using the r2SCANX orbitals. Y=100% includes the full Fock operator and strongly overestimates the band gaps.
@@ III Discussion
Hybrid XC functionals mix a fraction of exact HF exchange with DFA to improve the accuracy of the electronic structure, thereby correcting functional-driven errors, particularly the SIE. The amount of HF exchange remains a tunable parameter. Strategies proposed to tune the amount of HF exchange in the DFA include fitting molecular heat of formations to experimental data (Pople’s G1 dataset), for example, the B3PW91 and the B3LYP.66, 168 Appropriate amounts of HF can also be derived from first-principles perturbative theory calculations,69 as is the case for 25% HF exchange in PBE0 and HSE06.70, 71 Galli et al.63 demonstrated that the exact exchange fraction should vary like the inverse of the material’s dielectric constant. Scuseria et al.72, 73, 74, 75 proposed that the amount of HF can be determined as a function of the electron density at each point in space. We have shown that MiOjs’ oxidation energies generally improve when climbing the ladder from LSDA to PBE to r2SCAN, but further improvement is needed.
@@
III.A Electron Densities from Hartree-Fock, Hybrids, and DFT+U Approaches
HF also provides self-interaction-free, albeit uncorrelated, electronic charge densities. Electronic charge densities from HF or hybrid functionals can correct density-driven errors. We have demonstrated that HF (or hybrid functional) electronic charge densities can correct r2SCAN (or its hybrids) total energies for MiOjs by proposing a generalized r2SCANY@r2SCANX method. In r2SCAN@r2SCANX, we mix Yexact HF exchange in the DFA energy and set the fraction of exact HF exchange used to evaluate the charge density to X%. We have proposed optimal combinations of HF exchange in the charge density and functional definitions, such as r2SCAN10, r2SCAN10@r2SCAN, r2SCAN@r2SCAN50, and r2SCAN10@r2SCAN50, which reduce the mean absolute (and relative) errors of predicted electronic and magnetic properties, and oxidation energies of MiOjs.
@@ III.A Electron Densities from Hartree-Fock, Hybrids, and DFT+U Approaches
Kulik and collaborators demonstrated that the ability of DFT+U to localize electrons on transition metals can vary significantly from that of hybrid functionals; these latter tend to localize the minority spin density (of the transition metal) away from the metal, and towards oxygen atoms in transition metal complexes.169, 170 We observe the same behavior in MiOjs. In agreement with Kulik et al.,169, 170 we also show that a higher fraction of HF exchange in the XC functional increases the on-site magnetic moments on transition metal species in MiOjs, which is rationalized by the rise of majority spin in the d-manifolds. Considering these observations, we can qualitatively state that an increasing amount of HF in the DFA increases the ionic character of MiOjs. We quantify these behaviors for Fe2O3 Supplementary Fig. 5 by change in minority- and majority-spin electron numbers. In addition, Supplementary Fig. 6 shows a progressive increase in the electron number on the oxygen sites with increasing X in r2SCANX. This is preceded by a similar increase from LSDA to PBE to r2SCAN. The electron number on the transition-atom site decreases with increasing X. Exceptions to this trend are the MiOjs where strong charge disproportionation of the transition metal is favored, for example, Fe3O4 (Fe3+ and Fe2+), Mn3O4 (Mn3+ and Mn4+). As for the magnetic moments, the electron number on an atom is estimated by integrating the electron density over the PAW sphere.
In addition, while investigating transition metal complexes, Ref. 170 concluded that DFT+U charge densities might differ substantially from hybrid functionals. To test this observation, we compare r2SCAN@r2SCAN+U MiOjs band gaps and oxidation energies (Supplementary Fig. 7) to r2SCAN and r2SCAN10@r2SCAN50 results. We demonstrate that r2SCAN@r2SCAN+U oxidation energies and band gaps are comparable to r2SCAN.
@@ III.B Sources of Uncertainty and Error
The widely accepted standard for representing uncertainties in experimentally obtained oxidation enthalpies and other thermochemical data requires estimates of the 95% confidence intervals, which all thermochemical tables universally follow.149, 150, 151 Ruscic has argued that the MAE of calculated formation energies relative to experimental ones underestimates the error of the calculated values.171 For this reason, when benchmarking the accuracy of r2SCANY@r2SCANX (Fig. 3), we have also utilized the root mean square error.
@@
Reproducing errorless O2 binding energies is essential for accurate predictions of oxidation energies of MiOjs. LSDA and PBE drastically overbind the O2 molecule with large negative binding energies,29, 156, 157 $\sim$–2.2 eV/O2 and $\sim$–1 eV/O2, respectively, resulting in a systematic shift in predicted oxidation energies. With r2SCAN, this negative error drops in magnitude to a sizeable value of $\sim$–0.3 eV/O2. Applying a constant shift (of $\sim$0.3 eV/O2) to correct the spurious O2 overbinding in r2SCAN would further underestimate MiOj oxidation energies (see Fig. 2a), further exacerbating the error in r2SCAN predictions. Previously, similar corrections have been applied to PBE O2 and MiOj oxidation energies.29, 30 Notably, we have demonstrated that r2SCAN10@r2SCAN, r2SCAN10, and r2SCAN10@r2SCAN50 decrease the error in the O2 binding energy to values $\leq$ 0.03 eV/O2 in magnitude, and without requiring any fitting procedures.29, 30
@@
Since the errors of O2 binding energy are minimal with r2SCANY@r2SCANX methods, errors in the oxidation energies of MiOjs must reside in the prediction of the reduced and oxidized oxides. This is a bimodal distribution of r2SCAN oxidation energies (Fig. 2b), consisting of a low (near zero) peak for early transition metals, i.e., Ti, V, Cr, reactions, and higher errors for late transition metals, i.e., Mn, Fe, Co reactions. Notably, early transition metals (Ti, V, and Cr) involve a small number of d-electrons in their electronic configurations, probably reducing correlation effects, which become more prominent for late transition metals (Mn, Fe, and Co) with more electrons in d-shells. In addition, redox reactions involving exchanges between transition metals’ closed d-shell configurations to open d-shell configurations or vice versa, such as Ti4+(d0) $\rightarrow$ Ti3+ (d1), or Cu+ (d10) $\rightarrow$ Cu2+ (d9), appear to be in better agreement with experimental data. However, milder correlation effects in closed-shell transition metals make r2SCAN sufficient to describe the electronic structure of their oxides (for example, TiO2, V2O5, etc.).
@@
From these observations, we expect that the electron density by r2SCAN may be too delocalized to capture localization effects in late transition metals, i.e., Mn, Fe, Co, containing a larger number of d electrons. Better electronic charge densities as provided by r2SCAN@r2SCAN50, correct the electron number in the d manifolds, recovering a unimodal distribution (Fig. 2b), with a substantial decrease of oxidation energies errors for MiOjs with M = Mn, Fe, and Co. In addition, due to the functional-driven error, the average error of r2SCAN@r2SCAN50 is non-zero, resulting in an offset distribution around the zero-error line. However, in r2SCAN10@r2SCAN50 the functional-driven error reduces, and the unimodal distribution centers near the zero-error line.
@@
Magnetic moments are also typically impacted by SIE and are underestimated due to the overdelocalization of d(f) electrons on transition metal centers by GGA and meta-GGA functionals. As we climb the Jacob’s ladder172 of the XC functionals from LSDA (PW91) to GGA, and to meta-GGAs, the error in magnetic moments decreases accordingly, as observed from the functional MAE distributions of Fig. 5a and Fig. 5b. We observed that the error in magnetic moments drops slowly as the percentage of exact HF exchange in the orbital increases. Compared to r2SCAN, for r2SCAN50, the error in magnetic moment reduces by $\sim$15%, and for HF, the error reduces by $\sim$20%. In contrast, r2SCAN+U minimizes the error in magnetic moments much more mildly, $\sim$5%. Note that the variance in experimentally reported magnetic moments can be as high as +1.00 $\mu_{B}$. For example, magnetic moments in V3+ in V2O3 are reported to vary between 1.2 to 2.37 $\mu_{B}$ (Table 1),121, 122 which corresponds to variations larger than 100%. Despite this considerable uncertainty in experimentally reported magnetic moments, we observed a noticeable decrease in MAE of MiOjs predicted magnetic moments from r2SCAN to r2SCAN50.
@@
Band gaps of MiOjs are typically underestimated by standard DFT predictions, with absolute errors rapidly decreasing as LSDA (PW92) > GGA (PBE) $\gg$ meta-GGAs (r2SCAN and LAK) > hybrid meta-GGA (r2SCANX), see Fig. 6c and Fig. 6d. Using r2SCANY@r2SCANX, we demonstrated that errors in band gaps are minimized with approximately 5-10% exact exchange in the functional and the orbital definitions, which differs from the optimal HF fractions minimizing oxidation energy errors of MiOjs. We have shown that band gaps predicted by r2SCAN10@r2SCAN are more accurate than r2SCAN+U, with MAEs of 0.53 eV/O2 and 0.71 eV/O2, respectively.
III.C False Metals and Lattice Symmetry Breaking
@@
Materials possess many degrees of freedom embedded in their crystal structures. Deformation of octahedra (and cation polyhedra), octahedral tilting, and Jahn-Teller distortions all decrease the crystal symmetry in materials.24, 160 Additionally, specific magnetic orderings of transition metals in MiOjs, typically determined by neutron scattering and magnetic spectroscopies, further reduce crystal symmetries. On the one hand, standard structural techniques, such as X-ray and neutron diffraction, generally show less sensitivity to such distortions and defects. As such, reported crystal structures of materials may be inherently more symmetric than they appear. On the other hand, band-gap measurements are sensitive to symmetry-breaking motifs, defects in materials, and charge and magnetic orderings. Zunger and collaborators demonstrated that DFT of overly symmetrized MiOjs and other materials tends to close band gaps, predicting “false metals”.24, 158, 159, 160
@@
Starting from the highly symmetrized experimental structures, further optimized with r2SCAN, our r2SCAN band-gap predictions of Ti2O3,101 V2O3,103 and Fe3O4113 appear metallic (Table 1). Upon removing all possible symmetry layers (perturbing atomic positions from high-symmetry sites of large supercell models, imposing ground state magnetic orderings, removing intrinsic symmetry of wavefunctions, orbitals, and time reversal), Ti2O3, V2O3, and Fe3O4 remain metallic. However, the experimentally reported band gaps for Ti2O3, V2O3, and Fe3O4 are smaller than or equal to $\sim$0.2 eV (Table 1), which makes band-gap opening unlikely, even with substantial symmetry breaking as demonstrated by our calculations. We found no lattice symmetry breaking for Ti2O3 and V2O3, while we observed lattice symmetry breaking for Fe3O4. Note, there can be multiple symmetry breakings,24, 158 and we may not have found the one with the lowest energy. However, Supplementary Fig. 2 shows that (without lattice symmetry breaking) non-zero gaps in these materials arise in the self-consistent hybrid r2SCAN10.
III.D Computational Efficiency Considerations of r2SCANY@r2SCANX Approaches
It is relevant to comment on the computational costs faced by the r2SCANY@r2SCANX approach. We compared computational timings for r2SCAN10, r2SCAN10@r2SCAN, and r2SCAN on FeO (16 atoms per unit cell), Fe3O4 (14 atoms per unit cell), and Fe2O3 (10 atoms per unit cell). Since r2SCAN10 hybrid calculations were started from a converged r2SCAN calculation and r2SCAN10@r2SCAN used converged r2SCAN orbitals, we include the time to find the r2SCAN orbitals in the timings for r2SCAN10 and r2SCAN10@r2SCAN. The timing ratios of r2SCAN10 to r2SCAN were $\sim$77 for FeO, $\sim$562 for Fe3O4 (a false metal in r2SCAN and a small-gap insulator in r2SCAN10), and $\sim$120 for Fe2O3. The timing ratios of r2SCAN10@r2SCAN to r2SCAN were only $\sim$2.5 for FeO, $\sim$3.4 for Fe3O4, and $\sim$10 for Fe2O3.